OpenStudy (anonymous):

City A and City B had two different temperatures on a particular day. On that day, four times the temperature of City A was 8° C more than 3 times the temperature of City B. The temperature of City A minus twice the temperature of City B was –3° C. What was the temperature of City A and City B on that day? City A was 5° C, and City B was 4° C. City A was 3° C, and City B was –1° C. City A was 8° C, and City B was –3° C. City A was 5° C, and City B was –5° C.

OpenStudy (anonymous):

@Jasmiinnee_m

OpenStudy (anonymous):

@edenilson

OpenStudy (anonymous):

@AlexPinette

OpenStudy (anonymous):

The problem is giving you two equations here: 4a=3b+8 a-2b=-3 First isolate 'a' a=2b-3 Then plug that into equation one 4(2b-3)=3b+8 8b-12=3b+8 5b=20 b=4 Plug 'b' into easiest equation for finding 'a' a=2(4)-3 a=8-3 a=5 Answer: Temperature of city A=5 and city B=4

OpenStudy (anonymous):

thank alot can you help me alot